it takes 5 hours for a boat to go a distance of 15 miles against the current and 3 hours in the direction of the current. what is the speed of the boat in stil water and what is the speed of the current

Answer by mananth(16946)

boat speed =x mph
current speed =y mph
against current 5 hours
with current 3 hours

Distance against 15 miles distance with 15 miles
t=d/r against current -
15.00 / ( x - y )= 5.00
5.00 ( x - y ) = 3.00
5.00 x - 5.00 y = 15.00 ....................1

15.00 / ( x + y )= 3.00
3.00 ( x + y ) = 15.00
3.00 x + 3.00 y = 15.00 ...............2
Multiply (1) by 3.00
Multiply (2) by 5.00
we get
15.00 x + -15.00 y = 45.00
15.00 x + 15.00 y = 75.00
30.00 x = 120.00
/ 30.00
x = 4.00 mph

plug value of x in (1) y
5.00 x -5.00 y = 15.00
20.00 -5.00 -20.00 = 15.00
-5.00 y = 15.00
-5.00 y = -5.00 mph
y = 1.00
boat speed 4.00 mph
current speed 1.00 mph

m.ananth@hotmail.ca